# Genetic Data Governance in Crisis: Policy Recommendations for Safeguarding Privacy and Preventing Discrimination

### v1

@@ 6.3. Recommendation 3: A Genetic Data Regulation Framework
Recommendation: To eliminate leaky protections, we suggest a comprehensive and uniform regulatory framework that encompasses all usage domains whether genetic data is collected, analyzed, or used in any type of inference. Below, we build upon specific values inherent in privacy rights to ensure that responsibility is placed on the organizations that hold and analyze data, rather than the individuals from whom it was collected (Solove, 2023).
  1. (1)
@@
  2. (2)
    Test approval: After entity approval, we propose regulations for the inferential tests, specifically those that return results to either an individual or a medical practitioner for any actionable result. At a minimum, entities would be required to publicly release detailed white papers for each test that detail laboratory procedures, quality control steps, inferential models used, presentation of results, and any other necessary details that would allow one to recreate the analysis if given the same data.
  3. (3)
@@
    • Individuals should be able to request data removal. This would include specifically (1) the destruction of the physical sample, (2) deletion of data, (3) removal of any identifiers, and (4) that any removed data no longer influence the results of any downstream models.
@@
    • Research usage is common among DTC companies, academia, and hospitals, where data provided per a test can be pooled together as a dataset for research (both internally and via external collaborations). We suggest an opt-out strategy here for individuals, with a specification on whether they consent to their data becoming publicly available in any form (e.g. public datasets, open sourced GWAS weights, trained models, etc).
    • Incidental genetic discoveries can be possible in research use cases where data for a particular test was used for other tests. In these cases, where knowledge of a particular trait can also be harmful and uninformative to the individual, we suggest an opt-in strategy for an individual to blanket choose if they wish to hear any secondary discoveries, or forward any secondary discoveries to their medical practitioner, to be enacted on at their discretion.
  4. (4)
    Bankruptcy and Acquisition: As detailed by third-party data transfers, the owning of genetic data would also apply to companies who acquire any genetic data as part of assets through acquisitions or bankruptcy. Thus, all entities involved must already have approval to handle genetic data. In the rare event that there are no entities to handle the data, we suggest a similar protocol structure as nuclear waste (Doctorow, 2008), where all physical sample, data, and models are subsequently destroyed and cannot be recovered.
7. Conclusions and Open Challenges
@@
While our policy recommendations focus on enhancing user protections, we also identify three key challenges in regulating genetic data usage that fall outside the scope of our work and that would benefit from further research.
7.1. The relational theory of privacy.
7.1.1. Should genetic data submissions or inclusion into genetic datasets require consent from biological relatives?
@@
7.1.2. Who has ownership of children’s DNA?
Do parents have a “fiduciary duty” to protect their child’s genetic data and protect them from the system vulnerabilities that we have outlined? Part of this fiduciary duty would involve whether a genetic test should be taken in the first place, e.g., the difference between a medically-necessary test and a DTC genetic test19191923andMe’s policy for minors is as follows: individuals under 18 that give assent and have a parent/guardian’s permission can submit their sample (23andMe, 2024a). for some insights into the child’s ancestry and genealogical relatives. The relational aspect of genetic data also comes into play here; for example, should consent be required from both parents to sequence their child’s genome? See (Bala, 2023) for a thorough discussion of child’s genome ownership.
7.2. International genetic data transfers.
Our paper focuses on American regulatory and legislative bodies and American companies. In the DTC genetic testing space, individuals may be less likely to trust foreign companies: Nucleus Genomics [^20] https://mynucleus.com proudly states on their home page that customer genomes are sequenced in the U.S. and the data is never sent/stored overseas21212123andMe also genotypes samples in the U.S. (23andMe, 2024b), but this is not a selling point of theirs.. Should there be restrictions on genetic data transfer between countries? For example, one hypothetical scenario might involve a political refugee who has medically-necessary genetic testing done within the U.S, but because of leaky data transfer rules, genetic information is transferred to their home country where it might be used to identify family members now at risk for persecution.
7.3. The deployment of AI models for genetic data.
Our paper has largely shied away from discussions about genetic inference and AI for two reasons: (1) deployment of AI technologies for genetic inference is nascent and, most importantly, (2) risks and vulnerabilities exist even with current simple linear models for genetic inference. AI will, we believe, exacerbate the risks we have outlined as they will likely result in seemingly better trait prediction by modeling nonlinear interactions and taking advantage of gene-environment correlations(Yang et al., 2024). However, the uncertainties associated with the basic scientific inference (that we outline in Section 2) will be compounded because of the lack of interpretability and complexity of AI models.

### v2

@@ Introduction
Recommendation: We suggest a uniform regulatory framework that applies the same standards of collection, consent, and control to all data collectors in the genetic data ecosystem. In the spirit of the legitimacy challenge posed by Viljoen, we build upon specific values inherent in privacy rights to ensure that responsibility is placed on the organizations that hold and analyze data, rather than the individuals from whom it was collected, and the approach to consider data as a democratic medium (Solove 2022; Viljoen 2024).
  1. 1.
@@
  2. 2.
    Test approval: After entity approval, we propose regulations for the inferential tests, specifically those that return results to either an individual or a medical practitioner for any actionable result. At a minimum, entities would be required to publicly release detailed white papers for each test that detail laboratory procedures, quality control steps, inferential models used to validate the testing process for AI models, presentation of results, and any other necessary details that would allow one to recreate the analysis if given the same data.
  3. 3.
@@
    • Individuals should be able to request data removal. This would include specifically (1) destruction of the physical sample, (2) deletion of data, (3) removal of any identifiers, and (4) any removed data no longer influence the results of any downstream models.
@@
    • Research usage is common among DTC companies, academia, and hospitals, where data provided from individual tests can be pooled together as a dataset for research (both internally and via external collaborations). We suggest an opt-out strategy here for individuals, with a specification on whether they consent to their data becoming publicly available in any form (e.g. public datasets, open sourced GWAS weights, trained models, etc).
    • Incidental genetic discoveries can be possible in research use cases where data for a particular test was used for other tests. In these cases, where knowledge of a particular trait can also be harmful and uninformative to the individual, we suggest an opt-in strategy for an individual to blanket choose if they wish to hear any secondary discoveries, or forward any secondary discoveries to their medical practitioner.
  4. 4.
    Bankruptcy and Acquisition: As detailed by third-party data transfers, the owning of genetic data would also apply to companies who acquire any genetic data as part of assets through acquisitions or bankruptcy. Thus, all entities involved must already have approval to handle genetic data. In the rare event that there are no entities to handle the data, we suggest a similar protocol structure as nuclear waste, where all physical sample, data, and models are subsequently destroyed and cannot be recovered.
Conclusions and Open Challenges
@@
While our policy recommendations focus on enhancing individual protections, we also identify three key challenges in regulating genetic data usage that fall outside the scope of our work and that would benefit from further research.
The Relational Theory of Privacy
Should genetic data submissions require consent from biological relatives?
@@
Who has ownership of children’s DNA?
Do parents have a “fiduciary duty” to protect their child’s genetic data and protect them from the system vulnerabilities that we have outlined? Part of this fiduciary duty would involve whether a genetic test should be taken in the first place, e.g., the difference between a medically-necessary test and a DTC genetic test for insights into the child’s ancestry and genealogical relatives. The relational aspect of genetic data also comes into play; for example, should consent be required from both parents to sequence their child’s genome? See (Bala 2023) for a thorough discussion of child’s genome ownership.
International Genetic Data Transfers
Our paper focuses on U.S. regulatory and legislative bodies and U.S.-based companies. Should there be restrictions on genetic data transfer between countries? For example, one hypothetical scenario might involve a political refugee who has medically-necessary genetic testing done within the U.S, but because of leaky data transfer rules, genetic information is transferred to their home country where it might be used to identify family members now at risk for persecution.
Collective Rights for Genetic Data
Privacy scholars acknowledge the necessity of collective rights when building on the relational nature of genetic data. Government-controlled genetic databases present a fundamental challenge to democratic legitimacy, creating power imbalances that distort citizen-state relationships and potentially undermine collective political engagement. Safeguarding individual data-subject rights, while important, does not address the deeper issue: the lack of collective, democratic decision-making regarding whether, how, and for what purposes genetic data is created and used. Key scholars addressing collective rights include King, who demonstrates this phenomenon in DTC genetic testing (King 2019); Solow-Neiderman, who details how individual genetic submissions compromise collective privacy (Solow-Niederman 2021); Costello, who raises issues of group consent and responsibility (Costello 2022); and Viljoen, who argues for democratizing data for the public collective (Viljoen 2024).
While robust theoretical arguments support collective rights frameworks, their path to implementation is less clear. Our analysis bridges existing policy approaches with emerging theoretical frameworks, providing a scaffold for comprehensive genetic data governance. In an increasingly data-driven society, there is an urgent need to address major gaps in genetic data governance, which informs our process to develop a risk assessment framework and outline three concrete policy recommendations. This work represents not an endpoint but a call to action for policymakers, researchers, and ethicists to collaborate on transforming theoretical insights into practicable governance solutions that safeguard genetic privacy for present and future generations.